# Efficient Latent Variable Causal Discovery: Combining Score Search and Targeted Testing

### v2

@@ 9 Discussion and Conclusion
We have introduced a family of score-guided, mixed-strategy algorithms for causal discovery under latent variables and selection bias. BOSS-FCI and GRaSP-FCI extend the GFCI framework by substituting BOSS or GRaSP for FGES in the initial CPDAG search. LV-Lite provides a simple yet effective heuristic alternative that returns the PAG equivalence class of the BOSS DAG, scaling extremely well but lacking the ability to remove edges indicative of latent confounding. Finally, FCIT introduces a targeted-testing strategy that addresses a key statistical limitation of FCI-style methods.
A central innovation of FCIT is its integration of recursive path blocking into edge removal, in conjunction with discriminating paths, allowing separating sets to be identified using full path structure rather than adjacency subsets alone. Each update is checked for structural validity, guaranteeing that FCIT always produces a well-formed PAG. Our theoretical results establish the soundness of recursive blocking, the edge-minimality of the returned PAG, and the soundness of all orientations in the large-sample limit. Empirically, FCIT achieves consistently high adjacency and orientation precision, scales effectively to large models, and outperforms existing FCI-style methods in both accuracy and efficiency. Although not formally correct, LV-Lite remains a remarkably fast and accurate heuristic for very large problems.
@@
Finally, we would be remiss not to note an important opportunity for future software development relevant to all algorithms in this paper that begin with a BOSS initialization—LV-Lite, BOSS-FCI, and FCIT—for the linear Gaussian or linear non-Gaussian cases. Our implementations have focused on Java for portability and integration with the Tetrad platform, which already provides strong performance advantages over many alternatives. However, a recent C-language implementation, CBOSS, available in the causal-get GitHub package (https://github.com/bja43/causal-get, due to Bryan Andrews), achieves substantial speed improvements. [^9] The speed improvements are largely due to low-level linear algebra optimizations for calculating likelihoods. The practical extension to the linear non-Gaussian case is benchmarked in Andrews et al., [3]. For problems with 100 nodes and average degree 10—the “large” case profiled above—CBOSS completes in under a second with performance identical to our Java-based BOSS and scales efficiently to much larger problems. This opens the possibility of integrating CBOSS as an external module for the BOSS phase of our algorithms. In Java, such integration poses some engineering challenges, but a hybrid approach—running the BOSS stage in C and importing the resulting graph into Java for subsequent processing—would be straightforward. Alternatively, the remaining steps of our BOSS-based algorithms could themselves be reimplemented in C. Either direction represents a practical and worthwhile programming goal for future work.
In summary, FCIT provides a practical and statistically efficient alternative to existing FCI-style algorithms, consistently yielding well-formed PAGs with improved accuracy and scalability. Together with BOSS-FCI, GRaSP-FCI, and the heuristic LV-Lite, it expands the toolkit for causal discovery under latent variables—bridging the gap between correctness, scalability, and practical usability.
@@ A.1 GFCI Variants, Find Path to Target, Recursive Blocking
For FCIT, The crucial algorithm find_path_to_target is given in Algorithm 6, along with helper methods in Algorithm 3. This is used in the algorithm, block_paths_recursively, Algorithm 4, to construct sets B on the fly to block all paths between variables $X$ and $Y$. This uses the notion of reachability given in [7]. The optimization of the discriminating path rule R4 is given in Algorithm 7. Section A.2 gives the procedure we use to list discriminating and pre-discriminating paths.
As an intuition, the recursive blocking procedure underlying FCIT is closely related to the standard depth-first search algorithm for determining d-separation between variables $X$ and $Y$ given a conditioning set $\mathbf{Z}$ [7]. Its main innovation is to construct the conditioning set dynamically: as the algorithm explores paths from $X$ to $Y$, it incrementally builds a blocking set $\mathbf{B}$ by adding nodes that close open trails—“in a clever way.” [^10] We borrow this turn of phrase from Zhang, [17]. When the procedure halts and $X$ and $Y$ are not adjacent, it returns a final set $\mathbf{B_{final}}$ such that $\text{d-sep}(X,Y\mid\mathbf{B}_{final})$, or Null if no such separating set exists. The same logic applies to DAGs (directed acyclic graphs), CPDAGs (completed partially directed acyclic graphs), MAGs (mixed ancestral graphs), and PAGs (partial ancestral graphs). If $X$ and $Y$ are adjacent, the procedure instead returns a set that blocks all non-inducing paths between them—a property that FCIT directly exploits.
@@ B.6 Relationship to Zhang’s m-Separation
    A direct edge xyx*\!-\!*y has no interior vertices and therefore provides no opportunity to satisfy the blocking clauses by conditioning on a noncollider or leaving a collider unactivated. Hence, while clause (ii) (“all colliders are in or have descendants in BB”) holds trivially, this does not imply independence as long as $x$ and $y$ remain adjacent. To certify $m$-separation, both adjacency and path conditions must be satisfied. Zhang’s criterion [17] explicitly requires
@@ C.1 Zhang’s Final FCI Orientation Rules
    For completeness, we recap here the orientation rules (R0–R10) that Zhang [17] proves are both sound and complete for FCI in the presence of latent variables and selection bias. In these rules, $\alpha,\beta,\gamma,\theta$ denote generic vertices, and the symbol $\ast$” is a wildcard that can represent an arrowhead ($\rightarrow$), a tail (-), or a circle ($\circ$). Any rule that does not apply in a given configuration can be ignored.
@@ C.2 Optimizing R4 Using Recursive Blocking
    where each internal collider along the path is a parent of $Y$. In the standard approach [14], applying R4 often requires exploring many potential conditioning sets, which is time-consuming and can lead to inaccuracies in dense graphs.
@@ Appendix E Auxiliary Figures
    Figure 19: Adjacency Precision (AP) for 20-node graphs with average degree 2. Precision remains uniformly high across all algorithms (>0.93>0.93). BOSS-FCI GRaSP-FCI, and GFCI achieve near-perfect accuracy ($\approx 1.0$) throughout, with LV-Lite lagging behimd a bit. FCI is also slightly behind but still quite high. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 21: Arrowhead Precision (AHP) for 20-node graphs with average degree 2. LV-Lite lags at small samples, and FCI is behind overall; the other algorithms do quite well. FCI remains slightly lower but converges steadily. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 22: Arrowhead Recall (AHR) for 20-node graphs with average degree 2. Recall improves monotonically with NN for all algorithms, exceeding 0.7 at the largest sample size. FCI attains the highest overall recall, followed closely by FCIT, while LV-Lite and the GFCI-family algorithms show similar convergence. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 25: Tail Precision (TP) for 20-node graphs with average degree 2. Most algorithms are in a top group, with LV-Lite and FCI lagging behind. FCI trails slightly at larger sample sizes. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 28: Tail Path Precision for 20-node graphs with average degree 2. Precision varies strongly with NN; for large samples. LV-Lite is systematically behind. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 30: Proportion of well-formed PAGs for 20-node graphs with average degree 2. There is good PAG precision for this regime, with LV-Lite and FCIT returning perfect scores, as predicted. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 33: Adjacency Recall (AR) for 20-node graphs with average degree 6. LV-Lite and FCIT show the strongest gains with sample size, reaching $\approx 0.7$ at large NN. GFCI, BOSS-FCI, and FCI improve more modestly, and GRaSP-FCI remains lowest. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 34: Arrowhead Precision (AHP) for 20-node graphs with average degree 6. LV-Lite attains the highest precision and improves with NN; FCIT is next best. GFCI and BOSS-FCI remain around $0.7$, GRaSP-FCI is lower, and FCI lowest overall. There is a consistent dip at N=500N{=}500. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 35: Arrowhead Recall (AHR) for 20-node graphs with average degree 6. All curves rise with NN; LV-Lite and FCIT are highest at large NN ($\approx 0.6$), while the GFCI-family methods and FCI trail. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 36: Arrowhead Precision for Common Adjacencies (AHPC) for 20-node graphs with average degree 6. LV-Lite maintains the highest precision and improves with NN; FCIT also increases steadily. BOSS-FCI and GFCI remain around $0.72$, GRaSP-FCI is lower, and FCI lowest overall. There is a consistent dip at N=500N{=}500. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 37: Arrowhead Recall for Common Adjacencies (AHRC) for 20-node graphs with average degree 6. Recall increases with NN for all methods. At large $N$, BOSS-FCI is highest (near $0.93$), followed by GFCI and FCI; FCIT and GRaSP-FCI are close behind, and LV-Lite is slightly lower. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 39: Tail Recall (TR) for 20-node graphs with average degree 6. At the all sample sizes LV-Lite and FCIT have the highest scores, while the others falls short. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 40: Arrow Path Precision for 20-node graphs with average degree 6. LV-Lite and FCIT achieve the highest precision, increasing to $\approx 0.96$ at large NN. BOSS-FCI and GFCI cluster around $0.80$; GRaSP-FCI is lower and FCI lowest overall. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 42: Bidirected Latent Path Precision for 20-node graphs with average degree 6. There is weak performance on this statistic here. Precision peaks near intermediate NN and eases slightly at the largest NN. FCIT attains the highest precision at mid-range, with other algorithms trailing. LV-Lite orients no bidirected edges, yielding undefined precision, so it is not plotted. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 43: Proportion of well-formed PAGs for 20-node graphs with average degree 6. LV-Lite and FCIT consistently produce valid PAGs (=1.0=1.0), confirming structural correctness. Other algorithms fall off, sometimes sharply. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 44: Runtime (CPU milliseconds) for 20-node graphs with average degree 6. GRaSP-FCI incurs the highest cost, growing steeply at large NN ($\sim\!600$ ms). FCIT is moderate, while LV-Lite, LV-Lite, FCI, and GFCI remain consistently faster. Results are averaged over graphs with 0, 4, and 8 latent common causes.

### v3

@@ 9 Discussion and Conclusion
We have introduced a family of score-guided, mixed-strategy algorithms for causal discovery under latent variables and selection bias. BOSS-FCI and GRaSP-FCI extend the GFCI framework by substituting BOSS or GRaSP for FGES in the initial CPDAG search. LV-Dumb provides a simple yet effective heuristic alternative that returns the PAG equivalence class of the BOSS DAG, scaling extremely well but lacking the ability to remove edges indicative of latent confounding. Finally, FCIT introduces a targeted-testing strategy that addresses a key statistical limitation of FCI-style methods.
A central innovation of FCIT is its integration of recursive path blocking into edge removal, in conjunction with discriminating paths, allowing separating sets to be identified using full path structure rather than adjacency subsets alone. Each update is checked for structural validity, guaranteeing that FCIT always produces a well-formed PAG. Our theoretical results establish the soundness of recursive blocking, the edge-minimality of the returned PAG, and the soundness of all orientations in the large-sample limit. Empirically, FCIT achieves consistently high adjacency and orientation precision, scales effectively to large models, and outperforms existing FCI-style methods in both accuracy and efficiency. Although not formally correct, LV-Dumb remains a remarkably fast and accurate heuristic for very large problems.
@@
Finally, we would be remiss not to note an important opportunity for future software development relevant to all algorithms in this paper that begin with a BOSS initialization—LV-Dumb, BOSS-FCI, and FCIT—for the linear Gaussian or linear non-Gaussian cases. Our implementations have focused on Java for portability and integration with the Tetrad platform, which already provides strong performance advantages over many alternatives. However, a recent C-language implementation, CBOSS, available in the causal-get GitHub package (https://github.com/bja43/causal-get, due to Bryan Andrews), achieves substantial speed improvements. [^9] The speed improvements are largely due to low-level linear algebra optimizations for calculating likelihoods. The practical extension to the linear non-Gaussian case is benchmarked in andrews2023fast. For problems with 100 nodes and average degree 10—the “large” case profiled above—CBOSS completes in under a second with performance identical to our Java-based BOSS and scales efficiently to much larger problems. This opens the possibility of integrating CBOSS as an external module for the BOSS phase of our algorithms. In Java, such integration poses some engineering challenges, but a hybrid approach—running the BOSS stage in C and importing the resulting graph into Java for subsequent processing—would be straightforward. Alternatively, the remaining steps of our BOSS-based algorithms could themselves be reimplemented in C. Either direction represents a practical and worthwhile programming goal for future work.
In summary, FCIT provides a practical and statistically efficient alternative to existing FCI-style algorithms, consistently yielding well-formed PAGs with improved accuracy and scalability. Together with BOSS-FCI, GRaSP-FCI, and the heuristic LV-Dumb, it expands the toolkit for causal discovery under latent variables—bridging the gap between correctness, scalability, and practical usability.
@@ A.1 GFCI Variants, Find Path to Target, Recursive Blocking
For FCIT, The crucial algorithm find_path_to_target is given in Algorithm 6, along with helper methods in Algorithm 3. This is used in the algorithm, block_paths_recursively, Algorithm 4, to construct sets B on the fly to block all paths between variables $X$ and $Y$. This uses the notion of reachability given in [geiger1990d]. The optimization of the discriminating path rule R4 is given in Algorithm 7. Section A.2 gives the procedure we use to list discriminating and pre-discriminating paths.
As an intuition, the recursive blocking procedure underlying FCIT is closely related to the standard depth-first search algorithm for determining d-separation between variables $X$ and $Y$ given a conditioning set $\mathbf{Z}$ [geiger1990d]. Its main innovation is to construct the conditioning set dynamically: as the algorithm explores paths from $X$ to $Y$, it incrementally builds a blocking set $\mathbf{B}$ by adding nodes that close open trails—“in a clever way.” [^10] We borrow this turn of phrase from zhang2008completeness. When the procedure halts and $X$ and $Y$ are not adjacent, it returns a final set $\mathbf{B_{final}}$ such that $\text{d-sep}(X,Y\mid\mathbf{B}_{final})$, or Null if no such separating set exists. The same logic applies to DAGs (directed acyclic graphs), CPDAGs (completed partially directed acyclic graphs), MAGs (mixed ancestral graphs), and PAGs (partial ancestral graphs). If $X$ and $Y$ are adjacent, the procedure instead returns a set that blocks all non-inducing paths between them—a property that FCIT directly exploits.
@@ B.6 Relationship to Zhang’s m-Separation
    A direct edge xyx*\!-\!*y has no interior vertices and therefore provides no opportunity to satisfy the blocking clauses by conditioning on a noncollider or leaving a collider unactivated. Hence, while clause (ii) (“all colliders are in or have descendants in BB”) holds trivially, this does not imply independence as long as $x$ and $y$ remain adjacent. To certify $m$-separation, both adjacency and path conditions must be satisfied. Zhang’s criterion [zhang2008completeness] explicitly requires
@@ C.1 Zhang’s Final FCI Orientation Rules
    For completeness, we recap here the orientation rules (R0–R10) that Zhang [zhang2008completeness] proves are both sound and complete for FCI in the presence of latent variables and selection bias. In these rules, $\alpha,\beta,\gamma,\theta$ denote generic vertices, and the symbol $\ast$” is a wildcard that can represent an arrowhead ($\rightarrow$), a tail (-), or a circle ($\circ$). Any rule that does not apply in a given configuration can be ignored.
@@ C.2 Optimizing R4 Using Recursive Blocking
    where each internal collider along the path is a parent of $Y$. In the standard approach [spirtes2001causation], applying R4 often requires exploring many potential conditioning sets, which is time-consuming and can lead to inaccuracies in dense graphs.
@@ Appendix E Auxiliary Figures
    Figure 19: Adjacency Precision (AP) for 20-node graphs with average degree 2. Precision remains uniformly high across all algorithms (>0.93>0.93). BOSS-FCI GRaSP-FCI, and GFCI achieve near-perfect accuracy ($\approx 1.0$) throughout, with LV-Dumb lagging behimd a bit. FCI is also slightly behind but still quite high. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 21: Arrowhead Precision (AHP) for 20-node graphs with average degree 2. LV-Dumb lags at small samples, and FCI is behind overall; the other algorithms do quite well. FCI remains slightly lower but converges steadily. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 22: Arrowhead Recall (AHR) for 20-node graphs with average degree 2. Recall improves monotonically with NN for all algorithms, exceeding 0.7 at the largest sample size. FCI attains the highest overall recall, followed closely by FCIT, while LV-Dumb and the GFCI-family algorithms show similar convergence. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 25: Tail Precision (TP) for 20-node graphs with average degree 2. Most algorithms are in a top group, with LV-Dumb and FCI lagging behind. FCI trails slightly at larger sample sizes. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 28: Tail Path Precision for 20-node graphs with average degree 2. Precision varies strongly with NN; for large samples. LV-Dumb is systematically behind. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 30: Proportion of well-formed PAGs for 20-node graphs with average degree 2. There is good PAG precision for this regime, with LV-Dumb and FCIT returning perfect scores, as predicted. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 33: Adjacency Recall (AR) for 20-node graphs with average degree 6. LV-Dumb and FCIT show the strongest gains with sample size, reaching $\approx 0.7$ at large NN. GFCI, BOSS-FCI, and FCI improve more modestly, and GRaSP-FCI remains lowest. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 34: Arrowhead Precision (AHP) for 20-node graphs with average degree 6. LV-Dumb attains the highest precision and improves with NN; FCIT is next best. GFCI and BOSS-FCI remain around $0.7$, GRaSP-FCI is lower, and FCI lowest overall. There is a consistent dip at N=500N{=}500. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 35: Arrowhead Recall (AHR) for 20-node graphs with average degree 6. All curves rise with NN; LV-Dumb and FCIT are highest at large NN ($\approx 0.6$), while the GFCI-family methods and FCI trail. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 36: Arrowhead Precision for Common Adjacencies (AHPC) for 20-node graphs with average degree 6. LV-Dumb maintains the highest precision and improves with NN; FCIT also increases steadily. BOSS-FCI and GFCI remain around $0.72$, GRaSP-FCI is lower, and FCI lowest overall. There is a consistent dip at N=500N{=}500. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 37: Arrowhead Recall for Common Adjacencies (AHRC) for 20-node graphs with average degree 6. Recall increases with NN for all methods. At large $N$, BOSS-FCI is highest (near $0.93$), followed by GFCI and FCI; FCIT and GRaSP-FCI are close behind, and LV-Dumb is slightly lower. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 39: Tail Recall (TR) for 20-node graphs with average degree 6. At the all sample sizes LV-Dumb and FCIT have the highest scores, while the others falls short. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 40: Arrow Path Precision for 20-node graphs with average degree 6. LV-Dumb and FCIT achieve the highest precision, increasing to $\approx 0.96$ at large NN. BOSS-FCI and GFCI cluster around $0.80$; GRaSP-FCI is lower and FCI lowest overall. Results are averaged over graphs with 0, 4, and 8 latent common causes.
@@
    Figure 42: Bidirected Latent Path Precision for 20-node graphs with average degree 6. There is weak performance on this statistic here. Precision peaks near intermediate NN and eases slightly at the largest NN. FCIT attains the highest precision at mid-range, with other algorithms trailing. LV-Dumb orients no bidirected edges, yielding undefined precision, so it is not plotted. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 43: Proportion of well-formed PAGs for 20-node graphs with average degree 6. LV-Dumb and FCIT consistently produce valid PAGs (=1.0=1.0), confirming structural correctness. Other algorithms fall off, sometimes sharply. Results are averaged over graphs with 0, 4, and 8 latent common causes.
    Figure 44: Runtime (CPU milliseconds) for 20-node graphs with average degree 6. GRaSP-FCI incurs the highest cost, growing steeply at large NN ($\sim\!600$ ms). FCIT is moderate, while LV-Dumb, LV-Dumb, FCI, and GFCI remain consistently faster. Results are averaged over graphs with 0, 4, and 8 latent common causes.